(1 point) Animals and people that take in more energy than they expend will get fatter. Here are data on 12 rhesus monkeys: 6 lean monkeys (4% to 9% body fat) and 6 obese monkeys (13% to 14% body fat). The data report the energy expended in 24 hours (kilojoules per minute) and the lean body mass (kilograms, leaving out fat) for each monkey.

  Lean   Obese
  Body Mass   Energy   Body Mass   Energy
6.8 1.21 8.5 0.93
7.2 1.02 9.8 1.4
9.4 1.46 10.6 1.18
9.8 1.65 12.4 1.43
9.8 1.07 11.8 1.23
9.3 1.2 11.5 1.31

(a) What is the mean lean body mass of the lean monkeys?
ANSWER   x¯=   kilograms.

(b) What is the mean lean body mass of the obese monkeys?
ANSWER   x¯= kilograms.
Answer #1

Answer:

a)

To determine the mean lean body mass of the lean monkeys

Mean is given as the average of the given lean monkeys

i.e.,

Mean for lean monkeys = Sum of samples / Total samples

= (6.8 + 7.2 + 9.4 + 9.8 + 9.8 + 9.3) / 6

= 52.3 / 6

= 8.7167

Mean for lean monkeys = Xbar = 8.7167 kg

b)

To determine the mean lean body mass of the obese monkeys

Mean = Sum of samples / Total samples

= (8.5 + 9.8 + 10.6 + 12.4 + 11.8 + 11.5) / 6

= 64.6 / 6

= 10.7667

Mean for obese monkeys = Xbar = 10.7667 kg